Calculate the molar solubility of Zn(OH)2 given that Ksp at 25°C is 3.00×10-16.
Calculate the solubility (in grams per litre) of Zn(OH)2.
Answer #1

het molar Solubility is s. Zn(OH)2 -> Zint + 2. OH- s 25 Ksp = [Zx++) {ohr)? 3.00x1016 – S. (25)2 s=4.22*106 mol s=4.22x100x9